# Quantum Kernels for Symbolic Sequences Classification

### v1

@@ Abstract
This article presents a quantum computing approach to the design of similarity measures and kernels for classification of stochastic symbol time series. The similarity is estimated through a quantum generative model of the time series. We consider classification tasks where the class of each sequence depends on its future evolution. In this case a stochastic generative model provides natural notions of equivalence and distance between the sequences. The kernel functions are derived from the generative model, exploiting its information about the sequences evolution.We assume that the stochastic process generating the sequences is Markovian and model it by a Quantum Hidden Markov Model (QHMM). The model defines the generation of each sequence through a path of mixed quantum states in its Hilbert space. The observed symbols are emitted by application of measurement operators at each state. The generative model defines the feature space for the kernel. The kernel maps each sequence to the final state of its generation path. The Markovian assumption about the process and the fact that the quantum operations are contractive, guarantee that the similarity of the states implies (probabilistic) similarity of the distributions defined by the states and the processes originating from these states. This is the heuristic we use in order to propose this class of kernels for classification of sequences, based on their future behavior. The proposed approach is applied for classification of high frequency symbolic time series in the financial industry.
@@ 1 Introduction
Many machine learning algorithms for time series use various measures of similarity between examples to solve a wide range of problems, including classification, regression, density estimation, and clustering. The measures of similarity are critical part of the k-NN algorithms, feature-based algorithms, and kernel methods. Different similarity measures are intended to account for different types of similarity of the time series. One group of similarities are estimated directly in the original time series space. When the exact timing and synchronization of the data points are crucial the time series are compared synchronously, in lock-step, using an one-to-one alignment of the corresponding data points. Some common lock-step time series similarity measures include Euclidean Distance, Linear Time Warping (LTW)[1], Pearson Correlation Coefficient [2], Spearman Rank Correlation [3]. The synchronous methods are not applicable for series with different length and are unable to capture the local dependencies among adjacent positions within time series data. These problems can be resolved by using non-linear mappings between the time series which allow comparison of one to many points. This approach provides some degree of flexibility or ”warping” when aligning two time series. This flexibility makes it possible to find the optimal alignment between time series that may have variations in timing, speed, or phase. Examples of elastic similarity measures are the Longest Common Sub-sequence (LCSS) [4], Dynamic Time Warping (DTW) [5], Edit Distance [6], Move-Split-Merge (MSM) Distance [7], Piecewise Linear Approximation (PLA) Distance [8], Symbolic Aggregate Approximation (SAX) Distance [9], Elastic Ensemble Distance (EED) [10]. The synchronous and elastic similarities are estimated over the full time series and they are referred to as ”global similarities”. Other approaches measure the similarity with respect to local but representative patterns or ”shapelets” [11] within time series. For each time series, the similarity between the time series and each shapelet is calculated, and these values become distance features. Distance features can be defined by the global similarities measures as well. Any similarity measure creates a space - ”dissimilarity” feature spaces - where each time series is represented by its distance vector to the other time series.
Another group of very successful, similarity based methods - the kernel methods [12] - use a function which maps pairs of input series to the dot product of corresponding vectors in a high dimensional Hilbert space. The classical kernel functions calculate the dot product directly in the input space, without explicitly construction of vectors in the Hilbert space. If these functions are properly designed their outputs are measure of similarity between the input time series and between their high dimensional representations as well. The intention of the method is the high dimensional representations of the inputs to be linearly separated in the Hilbert space with respect to their class labels [13]. The kernel functions are possible due to the properties of the inner product structure within the Hilbert space [14]. Mercer’s theorem defines the necessary and sufficient conditions for a function to be a valid kernel function [13]. The kernel functions allow construction of linear discriminative algorithms for learning domains with non-linear decision boundaries. Examples of such algorithms are Support Vector Machines [15], Kernel Logistic and Ridge Regression [16], Kernel Discriminant Analysis [17], and many others. Kernel methods are used in unsupervised learning algorithms as well to capture non-linear structures in data: Kernel Principal Component Analysis [18, 19, 20], Non-linear Independent Component Analysis [21, 22], Kernel K-Means Clustering [23], Kernel Spectral Clustering [24], Kernel Density Estimation [25]. Numerous kernel functions have been introduced for input variables represented as fixed-length real vectors, including polynomial, inhomogeneous polynomial, Gaussian radial basis function, sigmoid kernel, and hyperbolic tangent kernels[12]. These kernels are not suitable in scenarios where the input variables consist of time series data with varying lengths. One simple solution to this problem is to align each vector to a common length. This can be accomplished by application of Linear Time Warping. A more sophisticated approach to the global alignment of sequences relies on optimized alignment and dynamic programming. The Dynamic Time Warping for example, optimally aligns two sequences by warping one to minimize dissimilarity between corresponding points. It computes local distances, constructs a distance matrix, and employs dynamic programming to find the optimal alignment, which is used to calculate the distance between the sequences. The distance is used to define the DTW kernel. If the goal is to estimate the significant local similarities then the method of local alignment score [26] could be used. The method uses dynamic programming and a scoring system to compute an optimal local alignment score, highlighting regions of strong similarity within the sequences. This score has been combined in [27] with a convolution kernel [28] to derive a local alignment kernel. For time series classification tasks with symbolic, noisy, or high dimensional data it is appropriate to use a kernel based on the Symbolic Aggregate Approximation (SAX) similarity measure [9].
The choice of a kernel type for particular machine learning task depends on the goal of the task and the specific characteristic of the process we are targeting. It would be valuable to have a systematic way to define the kernel class and parameters for a particular classification, regression or clustering task. Since a kernel function reflects an assumption about the metric relations between the time series, a possible approach is these metric relations to be defined from a probabilistic generative model which assigns likelihood to each time series. The underlying model’s distribution defines a feature space where the original data can be represented. The time series can be mapped to feature spaces of generation paths, likelihoods, likelihoods parameters, or likelihood gradients with respect to parameters. Generative model-based kernels provide a general way to measure similarity between data samples, especially when their underlying distribution is important. The method combines the power of generative models to capture data distributions with the flexibility of kernel methods for various machine learning tasks.
@@
There are multiple ways for implementing of generative model-based kernels. The P-kernels [23, 29] estimate the similarity or distance between two time series based on the joint probability of their inference over common generation paths of states. These kernels assume the existence of a probabilistic generative model, such as a Hidden Markov Model or Gaussian Mixture Model, which defines the probability distribution of the sequences. This distribution serves as a reproducing kernel space. In this approach, each time series is transformed into a feature vector of probabilities over all possible inference paths— sequences of states or sets of mixture components. The similarity of two series is then computed by taking the dot product of their respective feature vectors, with each component weighted by the probability of the corresponding inference path. Due to the feature vector containing probabilities for all possible inference paths, these kernels are also known as marginalisation kernels. In another generative approach, known as probability product kernel [30], each time series is mapped to a distribution within a parameterized class of distributions. The probabilistic models for the individual time series are identified using maximum likelihood estimation. The probability product kernel is defined as standard inner product between densities. If the inner product in the product kernel is replaced by the Kullback-Leibler divergence of the densities, the resulting kernel is the Kullback-Leibler kernel [31].
Since the vectors in the feature space are probability distributions, the distance between them can be estimated using common divergence measures like the Kullback-Leibler divergence, Jensen-Shannon divergence, or other statistical distance measures [31, 32]
@@
Another method for exploiting the knowledge captured by a probabilistic model to define a similarity metrics is to embed the time series in the gradient space of the generative model. This approach is known as Fisher Kernel [33]. The gradient vector of a series log-likelihood with respect to a model parameter describes how that parameter contributes to the generative process. These parameters can encompass transition and emission probabilities in the case of hidden Markov models, expectations, and variances in the context of Gaussian mixtures, and more. The gradient vector is known as Fisher score [34]. If two series have similar gradient vectors, it indicates the model generates them in a similar manner. Consequently, we consider the series being close to each other in terms of their generative processes. The Fisher kernel is defined as the inner product of the gradient vectors of any two time series.
In many discriminative learning tasks the class of a sequence depends on the distribution over its future evolutions. For example, the language models infer the most likely continuations of a sentence, the financial models estimate the probability a variable to be in a range in a future time, etc. In this article we study quantum models of discrete stochastic processes.
@@ 2.1 Stochastic Process Languages
The set of all finite sequences over the alphabet $\Sigma$, including the empty sequence $\epsilon$, is denoted by ${\Sigma}^{*}$. The set of all sequences with length exactly $t$ is denoted by ${\Sigma}^{t}$.
Any subset of ${\Sigma}^{*}$ is a language ${L}$ over the alphabet. The sequences belonging to a language are referred to as words.
The set of sequences originated by observations or measurements of the evolution of a discrete-time process is called the process language. It is easy to verify that if a word results from the observation of a process, then every one of its subwords has also been observed. Therefore, the process languages are subword-closed.
A stochastic process language $L$ is a process language together with a set
@@ 2.2 Probabilistic Generative Models
We will consider probabilistic generative models of stochastic process languages. These models assign probability to each sequence from $\Sigma^{*}$. Examples of such models are Variational Autoencoders (VAEs), Gaussian Mixture Models (GMMs), Hidden Markov Models (HMMs), Latent Dirichlet Allocation (LDA), Bayesian Networks, Boltzmann Machines, etc.
@@
where $S=\{s_{1}\ldots s_{n}\}$. The joint process
@@
is stationary and can be described by a linear model
@@
where at any moment in time $t$, the model is in a superposition (stochastic mixture) of its hidden states, described by a stochastic vector $x_{t}\in\mathbb{R}^{n}$. The component $x_{t}^{i}$ represents the probability of being in state $x_{t}=s_{i}$. The distribution of observable symbols at time at time $t$ is defined by the stochastic vector $y_{t}\in\mathbb{R}^{m}$ as the probability to observe $y_{t}=a_{i}$ is $y_{t}^{i}$. This model is known as (classical) hidden Markov model (HMM).
For every HMM $\mathbf{M}$ we can define a set of observable operators T as
$\textbf{T}=\{T_{a}:T_{a}=AB_{a},a\in\Sigma\}$ (7)
where $B_{a}=\operatorname{diag}(B[a,i]),i\in 1..n$ are diagonal matrices defining symbol observation probabilities for each state $s_{i}$ [35, 36].
Every component $T_{a}[i,j]$ of an observable operator $T_{a}$ defines the conditional probability to observe symbol $a$ when the state process evolves from state $s_{i}$ to $s_{j}$.
@@
Let $\textbf{a}={a_{1}}\ldots{a_{t}}$ be any sequence. We define an observable operator corresponding to a as follows:
@@
A Hidden Markov Model (HMM) $\mathbf{M}$ defines the probability of every finite observable sequence $\mathbf{a}=a_{1}\ldots a_{t}$ as:
@@
where 1 is the unit row vector with dimension $n$, and $\textbf{x}_{0}$ is an initial states distribution.
With every model $\mathbf{M}$ we associate a sequence function $f^{\mathbf{M}}:{\Sigma}^{*}\rightarrow\left[0,1\right]$ defined as:
@@ 2.3 Classification Problem for Stochastic Languages
We consider symbolic sequences classification problem for a stochastic process language $L$ and a finite set of class labels $C=\bigl{\{}c_{1},\cdots,c_{k}\bigr{\}}$. It is assumed that there exists an unknown probabilistic classification map $\mathcal{P}:L\rightarrow P[C]$ assigning to each sequence $\mathbf{y}\in L$ a distribution over the class labels. A convenient way to define the classification map is to use a mapping function
$p(\mathbf{y},c)=P[c|\mathbf{y}],\mathbf{y}\in L,c\in C$ (11)
that estimates the probability of each class label $c$ for a given data instance $\mathbf{y}$.
@@
where $Y=\bigl{\{}\mathbf{y}_{i}:\mathbf{y}_{i}\sim D^{L}_{|\mathbf{y}_{i}|},i\in\left[1,l\right]\bigr{\}}$ is a set of finite length sequences sampled from the distributions $D^{L}$ of $L$.
@@
The performance of the approximated mapping function $\hat{p}(\mathbf{y},c)$ is typically evaluated using probabilistic metrics such as log-likelihood, Brier score, or other probabilistic measures [???]. These metrics assess the model’s ability to predict the probability distribution over class labels and quantify the quality of the stochastic classification results.

### v2

@@ Abstract
This article presents a quantum computing approach to designing of similarity measures and kernels for classification of stochastic symbolic time series. In the area of machine learning, kernels are important components of various similarity-based classification, clustering, and regression algorithms. An effective strategy for devising problem-specific kernels is leveraging existing generative models of the example space. In this study we assume that a quantum generative model, known as quantum hidden Markov model (QHMM), describes the underlying distributions of the examples. The sequence structure and probability are determined by transitions within model’s density operator space. Consequently, the QHMM defines a mapping from the example space into the broader quantum space of density operators. Sequence similarity is evaluated using divergence measures such as trace and Bures distances between quantum states. We conducted extensive simulations to explore the relationship between the distribution of kernel-estimated similarity and the dimensionality of the QHMMs Hilbert space. As anticipated, a higher dimension of the Hilbert space corresponds to greater sequence distances and a more distinct separation of the examples. To empirically evaluate the performance of the kernels, we defined classification tasks based on a simplified generative model of directional price movement in the stock market. We implemented two widely-used kernel-based algorithms — support vector machines and k-nearest neighbors — using both classical and quantum kernels. Across all classification task scenarios, the quantum kernels consistently demonstrated superior performance compared to their classical counterparts.
@@ 1 Introduction
Many machine learning algorithms for time series use various measures of similarity between examples to solve a wide range of problems, including classification, regression, density estimation, and clustering. The measures of similarity play a critical role in the k-NN algorithms, feature-based algorithms, and kernel methods. Various measures are intended to capture different types of similarity within time series data. One category of similarities is estimated directly within the original time series space. In cases where precise timing and synchronization of data points are essential, time series are compared synchronously, in lock-step, using an one-to-one alignment of corresponding data points. Some common lock-step time series similarity measures include Euclidean distance, linear time warping (LTW)[1], Pearson correlation coefficient [2], Spearman rank correlation [3]. The synchronous methods are not applicable for series with different lengths and cannot capture the local dependencies among adjacent positions within time series data. These difficulties can be resolved by applying non-linear mappings between the time series which allow comparison of one to many points. This approach provides a degree of flexibility or “warping” when aligning two time series. Such flexibility makes it possible to find optimal alignment between time series that may exhibit variations in timing, speed, or phase. Examples of these elastic similarity measures include the longest common sub-sequence (LCSS) [4], dynamic time warping (DTW) [5], edit distance [6], move-split-merge (MSM) distance [7], piecewise linear approximation (PLA) distance [8], symbolic aggregate approximation (SAX) distance [9], elastic ensemble distance (EED) [10]. The synchronous and elastic similarities are estimated over the entire time series and are often referred to as “global similarities”. Alternatively, some approaches measure similarity with respect to local but representative patterns known as “shapelets” [11] within time series. For each time series, the similarity between the time series and each shapelet in a “dictionary” is calculated, resulting in distance features. Distance features may also be defined by global similarity measures. Any similarity measure is associated with a “dissimilarity” feature space, where each time series is represented by its distance vector to the other time series.
Another group of highly successful, similarity-based methods, known as kernel methods [12] utilize a function that estimates the similarity between examples by computing the dot product of their images in a higher-dimensional feature space. When the kernel functions satisfy Mercer’s theorem [13, 14] they can calculate the dot product directly in the example space, without explicitly construction of feature vectors [15]. The objective of this method is to ensure that the high-dimensional feature representations of the examples are linearly separable with respect to their class labels [14]. Kernel methods enable the construction of linear discriminative algorithms for learning domains with non-linear decision boundaries. Examples of such algorithms include support vector machines [15], kernel logistic and ridge regression [16], and kernel discriminant analysis [17], among others. Kernel methods are also used in unsupervised learning algorithms to capture non-linear structures in data. These include kernel principal component analysis [18, 19, 20], non-linear independent component analysis [21, 22], kernel K-means clustering [23], kernel spectral clustering [24], and kernel density estimation [25].
Most general kernel functions, such as polynomial functions, Gaussian radial basis functions, sigmoid kernels, and hyperbolic tangent functions, rely on simple pairwise distances, such as the Euclidean distance, between data points in the example space. However, these kernels may not be suitable when the example space contains time series with varying lengths. One direct solution to this problem is to utilize elastic similarity measures as demonstrated in the case of DTW kernels [26, 27]. If the goal is to estimate significant local similarities the method of local alignment score [28] can be used to define the kernel. This approach has been combined in [29] with a convolution kernel [30] to derive a local alignment kernel. Finally, for time series classification tasks involving symbolic, noisy, or high-dimensional data, it is appropriate to use a kernel based on the symbolic aggregate approximation (SAX) similarity measure [9].
The selection of kernel type for a specific machine learning task depends on the nature of the task and the specific characteristics of the process under consideration. It is important to establish a systematic method to define the kernel class and parameters for any particular learning scenario.
@@
One approach to kernel design involves leveraging information from an existing, unsupervised, generative model of the examples domain. Such models define underlying distributions within the examples domain. The underlying distributions define feature spaces where the examples are represented. There are multiple ways for implementing of generative model-based kernels. The P-kernels [23, 31] estimate the similarity between two time series based on the joint probability of their inference state sequences, also known as inference paths. These kernels assume the existence of a probabilistic generative model, such as a hidden Markov model or Gaussian mixture model. For a time series, the model defines probabilities of all its inference paths — sequences of model states or mixture components. This distribution serves as the kernel’s feature space. In this approach, each time series is transformed into a feature vector of probabilities over all possible inference paths. The similarity between two series is then computed by taking the dot product of their respective feature vectors, with each component weighted by the probability of the corresponding inference path. Because the feature vector contains probabilities for all possible inference paths, these kernels are also referred to as marginalisation kernels.
In another generative approach, known as probability product kernel [32], each time series is mapped to a distribution within a parameterized class of distributions. The probabilistic models for the individual time series are identified using maximum likelihood estimation. The probability product kernel is defined as standard inner product between densities. Since the vectors in the feature space are probability distributions, the distance between them can also be estimated using common divergence measures like the Kullback-Leibler divergence, Jensen-Shannon divergence, or other statistical distance measures [33, 34]. It is important to note, that these kernel functions usually do not satisfy the Mercer’s theorem, which can lead to various issues in kernel-based learning algorithms.
@@
One of the first methods for exploiting the knowledge captured by a probabilistic model to define a similarity metrics is to embed the time series in the gradient space of the generative model. This approach is known as Fisher kernel [35]. The gradient vectors of a series log-likelihood with respect to a model parameters describe how that parameters contribute to the generative process. These parameters can encompass transition and emission probabilities in the case of hidden Markov models, expectations, and variances in the context of Gaussian mixtures, and more. The gradient vector is known as Fisher score [36]. If two series have similar gradient vectors, it indicates the model generates them in a similar manner. Consequently, we consider the series being close to each other in terms of their generative processes. The Fisher kernel is defined as the inner product of the gradient vectors of any two time series.
In this article, we investigate similarity measures and their associated kernels for symbolic time series sampled from an underlying stochastic process language. We assume the existence of a Markovian quantum generative model of the language [37], which defines two classes of probability distributions: one over the observed sequences and the other over their inference paths of quantum states represented by quantum density operators. The feature spaces of our kernels are the spaces of quantum density operators. We estimate similarity in these spaces using divergence measures such as trace and Bures distances.
The article’s organization is as follows: section 2 provides a brief introduction to concepts, definitions, and features related to stochastic process languages and the corresponding generative models. We formalize classification problems for symbolic time series based on the concept of a classification map. Depending on the domain of the classification map, we define two types of classification tasks: structural and predictive. We discuss the definitions and requirements of the kernel functions utilized in classification problems. In section 3 we define a quantum generative model — quantum hidden Markov model (QHMM) — which serves as the foundation of our study. The model is defined as a completely positive trace-preserving (CPTP) map, enabling comprehensive definitions of symbolic sequence probabilities in the example space and generative state sequences in quantum density space. Subsequently, we discuss an unitary physical implementation of the model introduced in [37]. In section 4, we formally specify quantum predictive and structural kernels based on the introduced QHMM. We explain the intuition behind these kernels and establish some important probabilistic properties. Section 5 presents the conducted simulations and empirical investigations regarding the impact of the QHMM’s Hilbert space dimension and the distance distributions in the context of classification tasks. Here, we empirically compare the performance of the introduced quantum kernels with their classical counterparts. The discussion of the physical implementation of the proposed kernels in the quantum circuits computing model is presented in section 6. The contributions of the article, its potential impact on the field, and the future research directions are presented in section 7.
@@ 2.1 Stochastic Process Languages
The set of all finite sequences over the alphabet $\Sigma$, including the empty sequence $\epsilon$, is denoted by ${\Sigma}^{*}$. The set of all sequences with length exactly $t$ is denoted by ${\Sigma}^{t}$. Any subset of ${\Sigma}^{*}$ is a language ${L}$ over the alphabet. The sequences belonging to a language are referred to as words.
The set of sequences resulting from observations or measurements of the evolution of a discrete-time process is called a process language. In this context, the index $t$ is non-negative and is interpreted as time with the process sequences being considered as time series. It is straightforward to verify that if a word results from the observation of a process, then every one of its subwords has also been observed. Therefore, the process languages are subword-closed.
A stochastic process language $L$ is defined as a process language along with a set
@@ 2.2 Probabilistic Generative Models
We will consider probabilistic generative models for the stochastic process languages. These models describe the underlying distribution of the language assigning probability to each sequence from $\Sigma^{*}$. Examples of such models are variational autoencoders (VAEs), Gaussian mixture models (GMMs), hidden Markov models (HMMs), latent Dirichlet allocation (LDA), Bayesian networks, Boltzmann machines, etc [38].
@@
where $S=\{s_{1}\ldots s_{n}\}$ are the process states.
The joint process
@@
is assumed to be stationary and is described by a linear model
@@
where $A$ is a row-stochastic transition matrix and $B$ is column-stochastic emission matrix.
At any moment in time $t$, the model is in a superposition (a stochastic mixture) of its hidden states, described by a stochastic vector $x_{t}\in\mathbb{R}^{n}$. The state component $x_{t}^{i}$ represents the probability of the process being in state $x_{t}=s_{i}$. The distribution of observable symbols at time $t$ is defined by the stochastic vector $y_{t}\in\mathbb{R}^{m}$ as the probability to observe $y_{t}=a_{i}$ is denoted by $y_{t}^{i}$. This model is known as a (classical) hidden Markov model (HMM).
For every HMM $\mathbf{M}$ we can define a set of observable operators T as follows
$\textbf{T}=\{T_{a}:T_{a}=AB_{a},a\in\Sigma\},$ (7)
where $B_{a}=\operatorname{diag}(B[a,i],i\in[1.n])$ are diagonal matrices and $B[a,i]$ represents the emission probability of observing symbol $a$ in state $s_{i}$  [39, 40].
Every component $T_{a}[i,j]$ of an observable operator $T_{a}$ defines the conditional probability to observe symbol $a$ when the process state evolves from state $s_{i}$ to $s_{j}$.
@@
For any sequence $\textbf{a}={a_{1}}\ldots{a_{t}}$ we also can define an observable operator as follows:
@@
A HMM $\mathbf{M}$ defines probability of every observable sequence $\mathbf{a}=a_{1}\ldots a_{t}$ as:
@@
where 1 is the unit row vector with dimension $n$, and $\textbf{x}_{0}$ is the initial state distribution of the model. With every model $\mathbf{M}$ we associate a sequence function $f^{\mathbf{M}}:{\Sigma}^{*}\rightarrow\left[0,1\right]$ defined as:
@@ 2.3 Classification Problem for Stochastic Languages
We consider symbolic time series classification problem for a stochastic process language $L$ and a finite set of class labels $C=\bigl{\{}c_{1},\cdots,c_{k}\bigr{\}}$. It is assumed that there exists an unknown probabilistic classification map $\mathcal{P}:L\rightarrow P[C]$ assigning to each sequence $\mathbf{y}\in L$ a distribution over the class labels. We consider two types of probabilistic classification maps. The first classification map depends solely on the structure of a given sequence $\mathbf{y}$ and is defined by a mapping function $p_{s}$ estimating the probability of each class label $c$ for $\mathbf{y}$:
$p_{s}(\mathbf{y},c)=P[c|\mathbf{y}],\mathbf{y}\in L,c\in C.$ (11)
We refer to classification problems with mapping function (11) as structural. For such problems, the class of a sequence is directly associated with its structure. Examples of structural problems include the classification of protein sequences into functional families [41], detection of fraudulent financial transactions [42], and monitoring for threats in the cybersecurity domain [43] ], to name a few.
In another scenario the classification map depends on the probabilities and structure of the future evolution or continuations with length $k$ of a sequence $\mathbf{y}$ and is defined as:
$p_{p}(\mathbf{y},k,c)=\sum_{\mathbf{z}\in\Sigma^{k}}(P[\mathbf{z}|\mathbf{y}]P[c|\mathbf{z}]),\mathbf{y}\in L,c\in C,$ (12)
where $P[\mathbf{z}|\mathbf{y}]$ is the probability to observe $\mathbf{z}$ as continuation of $\mathbf{y}$. These types of classification problems are known as predictive. Predictive classification problems find applications in various domains such as [44], stock market prediction [45], health monitoring [46], energy load forecasting [47], and traffic flow prediction [48].
@@
where $Y=\bigl{\{}\mathbf{y}:\mathbf{y}\sim D^{L}_{i},i\in\left[1,l\right]\bigr{\}}$ is a set of finite length sequences sampled from the distributions $D^{L}_{i}$ of $L$.
@@
The performance of the approximated mapping function $\hat{p}(\mathbf{y},c)$ is typically evaluated using probabilistic metrics such as log-likelihood, Brier score, or other probabilistic measures. These metrics assess the model’s ability to predict the probability distribution over class labels and quantify the quality of the stochastic classification results.